The length of an aluminum pipe at 24 Celsius is 47.0cm. What is it’s length at 66 Celsius? State your answer in cm to nearest 0.001 cm, follow the standard rounding rules in stating answer. Take the thermal expansion coefficient of aluminum to be 24*10^-6/K, and take the absolute zero to be -273 Celsius.


1
Expert's answer
2021-10-02T14:56:05-0400

"l_l=l_0(1+\alpha t)\to l_0=l_1/(1+\alpha t)=0.47/(1+24\cdot10^{-6})=46.999\ (cm)"


"l_2=46.999\cdot(1+24\cdot10^{-6}\cdot 66)=47.073\ (cm)" . Answer